# Framing global structural identifiability in terms of parameter symmetries

### v1

@@ 3 Results
Given the notion of parameter invariants, we proceed by formulating the notion of global structural identifiability in terms of parameter symmetries.
3.2 Global structural identifiability defined in terms of universal parameter invariants
We now present our main result expressing global structural identifiability in terms of universal parameter invariants.
Theo. 1 (A parameter is globally structurally identifiable if it is a universal parameter invariant).
Let $\mathbf{y}(t,\boldsymbol{\theta})$ be a solution of the system of output ODEs in Eq. (2). A parameter $\theta_{\ell}\in\boldsymbol{\theta}$, ${\ell}\in\{1,\ldots,p\}$ is (globally) structurally identifiable if and only if $\theta_{\ell}$ is a universal parameter invariant.
@@ 3.2 Global structural identifiability defined in terms of universal parameter invariants
$\Rightarrow$” We need to prove that a structurally identifiable parameter $\theta_{\ell}$ satisfying the implication $\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\boldsymbol{\theta}^{\star})\implies{\theta}_{\ell}=\theta_{\ell}^{\star}$ for almost all values $\boldsymbol{\theta}^{\star}$ is itself a universal parameter invariant. Take a particular parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}:\boldsymbol{\theta}\mapsto\hat{\boldsymbol{\theta}}(\varepsilon)$. This parameter symmetry acts continuously on the parameters, and it leaves both the independent time variable $t$ and the dependent output variables $\mathbf{y}$ invariant according to Eq. (18). Therefore, we must have that
Using the previously mentioned implication on Eq. (26), since the parameter $\theta_{\ell}$ is globally structurally identifiable it follows that
@@
$\theta_{\ell}=\hat{\theta}_{\ell}(\varepsilon)\quad\forall\varepsilon\in\mathbb{R},$ (27)
implying that $\theta_{\ell}$ is a parameter invariant of $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}$ according to Eq. (25). Moreover, the same argument holds for all parameter symmetries and thus $\theta_{\ell}$ is a universal parameter invariant.
$\Leftarrow$” We need to prove that a universal parameter invariant $\theta_{\ell}$ is also structurally identifiable, and in this case we argue by contradiction. Let $\theta_{\ell}$ be a universal parameter invariant that is not structurally identifiable. This implies that for some parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}:\boldsymbol{\theta}\mapsto\hat{\boldsymbol{\theta}}(\varepsilon)$ and some transformation parameter $\varepsilon^{\star}$ the implication $\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\boldsymbol{\theta}^{\star})\implies{\theta}_{\ell}=\theta_{\ell}^{\star}$ does not hold for the particular value $\boldsymbol{\theta}^{\star}=\hat{\boldsymbol{\theta}}(\varepsilon^{\star})$. Specifically, we have that $\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\hat{\boldsymbol{\theta}}(\varepsilon^{\star}))$ while simultaneously $\theta_{\ell}\neq\hat{\theta}_{\ell}(\varepsilon^{\star})$. But since $\theta_{\ell}$ is a universal parameter invariant it satisfies Eq. (27) for all parameter symmetries, and hence we have a contradiction. Thus, there cannot exist such a parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ and such a transformation parameter $\varepsilon^{\star}$ and $\theta_{\ell}$ is structurally identifiable. ∎
@@
We say that a model is structurally identifiable if all parameters are globally structurally identifiable. In light of Thm. 1, a model is structurally identifiable if all parameters are universal parameter invariants. Given our previous discussion about the number of parameter invariants of parameter symmetries, an equivalent formulation is that a model is structurally identifiable if the only parameter symmetry of its input-output system is the trivial parameter symmetry $\Gamma_{0}^{\boldsymbol{\theta}}:\boldsymbol{\theta}\mapsto\boldsymbol{\theta}$.
@@
When a model is structurally unidentifiable, it is of interest to find the structurally identifiable parameter groupings or parameter quantities. Assume that the parameter symmetries $\Gamma_{\varepsilon}$ of interest have $\tilde{p}\leq{p}$ universal parameter invariants denoted by $I_{k}$, $k\in\{1,\ldots,\tilde{p}\}$ which are collected in a vector $\vec{I}_{\boldsymbol{\theta}}$. Crucially, we can always re-parametrise outputs $\mathbf{y}(t,\boldsymbol{\theta})$ in terms of these universal parameter invariants giving us $\mathbf{y}(t,\vec{I}_{\boldsymbol{\theta}})$, and then apply Thm. 1 on the re-parametrised outputs to give (Cor. 1).
@@
Cor. 1 (Global structural identifiability in terms of universal parameter invariants).
@@
The (globally) structurally identifiable parameter quantities of the system of output ODEs in Eq. (2) are given by its universal parameter invariants.
Subsequently, we use a toy example to illustrate that calculating the universal parameter invariants provides information that is consistent with the findings of the standard differential algebra approach.
3.2.1 Analysing the structural identifiability of a toy model using parameter symmetries
Let $u(t)$ and $v(t)$ denote the concentrations of two chemical species depending on time $t$ which satisfy two decoupled decay ODEs
$\dot{u}=\kappa_{1}-\lambda{u},\quad\dot{v}=\kappa_{2}-\lambda{v},$ (28)
where both species decay at rate $\lambda>0$ and are synthesised at rates rates $\kappa_{1}$ and $\kappa_{2}$, respectively. Furthermore, assume that we observe the total concentration $y(t)=u(t)+v(t)$, yielding the following model for the output $y$:
$\dot{y}=(\kappa_{1}+\kappa_{2})-\lambda{y}.$ (29)
We first apply the standard differential algebra approach for analysing the structural identifiability of this model. To this end, we extract the coefficients in front of $\{\dot{y},y,1\}$ resulting in the set $\{1,\lambda,\kappa_{1}+\kappa_{2}\}$. Clearly, the set of identifiable parameter quantities is given by $\{\lambda,\kappa_{1}+\kappa_{2}\}$ implying that the decay rate $\lambda$ is (globally) structurally identifiable. On the other hand, the synthesis rates $\kappa_{1}$ and $\kappa_{2}$ are individually unidentifiable whereas the sum $\kappa_{1}+\kappa_{2}$ is structurally identifiable.
@@ 3.2.1 Analysing the structural identifiability of a toy model using parameter symmetries
Next, we consider the symmetry-based approach for elucidating structural identifiability which entails finding universal parameter invariants of the output ODE in Eq. (29). To this end, we look for parameter symmetries $\Gamma_{\varepsilon}^{\theta}$ of the output ODE in Eq. (29) with the following structure
$\Gamma_{\varepsilon}^{\theta}:(\kappa_{1},\kappa_{2},\lambda)\mapsto(\hat{\kappa}_{1}(\kappa_{1},\kappa_{2},\lambda,\varepsilon),\hat{\kappa}_{2}(\kappa_{1},\kappa_{2},\lambda,\varepsilon),\hat{\lambda}(\kappa_{1},\kappa_{2},\lambda,\varepsilon)).$ (30)
@@
We denote the generating vector field of the parameter symmetry $\Gamma_{\varepsilon}^{\theta}$ in Eq. (30) by
$X_{\theta}=\chi_{\kappa_{1}}(\kappa_{1},\kappa_{2},\lambda)\partial_{\kappa_{1}}+\chi_{\kappa_{2}}(\kappa_{1},\kappa_{2},\lambda)\partial_{\kappa_{2}}+\chi_{\lambda}(\kappa_{1},\kappa_{2},\lambda)\partial_{\lambda}.$ (31)
@@
Given this vector field, we consider the following linearised symmetry condition of our toy model
@@
$X_{\theta}(\dot{y}-[(\kappa_{1}+\kappa_{2})-\lambda{y})]=0\quad\textrm{whenever}\quad\dot{y}-[(\kappa_{1}+\kappa_{2})-\lambda{y}]=0,$ (32)
@@
which states that the solution manifold is invariant under transformations by the parameter symmetry $\Gamma_{\varepsilon}^{\theta}$ in Eq. (30). Carrying out the differentiation on the left-hand side yields the following equivalent equation
@@
$X_{\theta}=\chi_{\kappa_{1}}(\kappa_{1},\kappa_{2},\lambda)(\partial_{\kappa_{1}}-\partial_{\kappa_{2}}),$ (35)
@@
$X_{\theta}(I(\kappa_{1},\kappa_{2},\lambda))=\chi_{\kappa_{1}}(\kappa_{1},\kappa_{2},\lambda)\left(\frac{\partial I}{\partial\kappa_{1}}-\frac{\partial I}{\partial\kappa_{2}}\right)=0.$ (36)
@@
where $I_{2}$ is an arbitrary integration constant. Since $I_{2}$ is a first integral of Eq. (42), this implies that it is also another differential invariant. In total, this implies that the two universal parameter invariants of our parameter symmetry $\Gamma_{\varepsilon}^{\theta}$ in Eq. (30) are given by
@@
$\Gamma_{\varepsilon}^{\theta}:(\kappa_{1},\kappa_{2},\lambda)\mapsto(\kappa_{1}+\varepsilon,\kappa_{2}-\varepsilon,\lambda),$ (45)

### v3

@@ 3 Local structural identifiability and parameter symmetries
Given the notion of parameter invariants, we proceed by formulating the notion of local structural identifiability in terms of parameter symmetries.
3.2 Local structural identifiability defined in terms of universal parameter invariants
One of the key properties of pointwise Lie symmetries of one parameter, in general, and parameter symmetries, in particular, is that these transformations are defined by single-variable, real-valued analytic functions. Based on this property, we now present our main result expressing structural identifiability in terms of universal parameter invariants.
Theo. 1 (Local structural identifiability of individual parameters in terms of universal parameter invariants).
Let $\mathbf{y}(t,\boldsymbol{\theta})$ be a solution of the system of output ODEs in Eq. (4). A parameter $\theta_{\ell}\in\boldsymbol{\theta}$, ${\ell}\in\{1,\ldots,p\}$ is locally structurally identifiable if and only if it is a universal parameter invariant.
@@ 3.2 Local structural identifiability defined in terms of universal parameter invariants
$\Longrightarrow$” We need to prove that a locally structurally identifiable parameter $\theta_{\ell}$ is also a universal parameter invariant. To this end, take a radius $\rho>0$ and a centre $\boldsymbol{\theta}_{C}\in\mathbb{R}^{p}$ defining a ball $\mathcal{B}_{\rho}\left(\boldsymbol{\theta}_{C}\right)$ in which the parameters are identifiable, which means that the following holds:
$\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\boldsymbol{\theta}^{\star})\quad\forall{t}\in\mathbb{R}\implies{\theta}_{\ell}=\theta_{\ell}^{\star}\quad\quad\forall\boldsymbol{\theta},\boldsymbol{\theta}^{\star}\in\mathcal{B}_{\rho}\left(\boldsymbol{\theta}_{C}\right)\,.$ (17)
Take a parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$. Since this parameter symmetry acts infinitesimally and pointwise, starting from any parameter $\boldsymbol{\theta}\in\mathcal{B}_{\rho}(\boldsymbol{\theta}_{C})$ we can always find transformation parameters $\varepsilon_{1},\varepsilon_{2}\in\mathbb{R}$ where $\varepsilon_{1}<\varepsilon_{2}$ such that these parameters form an interval $\varepsilon_{1},\varepsilon_{2}]$ for which the corresponding transformed parameters $\hat{\boldsymbol{\theta}}(\varepsilon)$ lie in the ball where the parameters are locally identifiable, which means that $\hat{\boldsymbol{\theta}}(\varepsilon)\in\mathcal{B}_{\rho}\left(\boldsymbol{\theta}_{C}\right)$ for all $\varepsilon\in[\varepsilon_{1},\varepsilon_{2}]$. Also, the independent variable (i.e. time, $t$) and the dependent variables (i.e. the outputs, $\mathbf{y}$) are invariant under transformations by the parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$, and thus the following conservation property holds as well: $\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\hat{\boldsymbol{\theta}}(\varepsilon))$ for all $t,\varepsilon$$\in\mathbb{R}$. Consequently, it follows from Eq. (17) that
@@
$\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\hat{\boldsymbol{\theta}}(\varepsilon))\quad\forall{t}\in\mathbb{R}\implies{\theta}_{\ell}=\hat{\theta}_{\ell}(\varepsilon)\quad\quad\forall\varepsilon\in[\varepsilon_{1},\varepsilon_{2}]\,$,
whenever $\theta_{\ell}$ is locally structurally identifiable. This means that $\hat{\theta}_{\ell}(\varepsilon)$ is constant in the subdomain $\varepsilon_{1},\varepsilon_{2}]$. Since $\hat{\theta}_{\ell}(\varepsilon)$ is a single variable real-valued analytic function, it must therefore be constant throughout its entire domain, which means that $\hat{\theta}_{\ell}(\varepsilon)={\theta}_{\ell}\,\,\forall\varepsilon\in\mathbb{R}$, and thus $\theta_{\ell}$ is a parameter invariant of the parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$. The same arguments hold for all parameter symmetries, and thus $\theta_{\ell}$ is a universal parameter invariant.
[…]
$\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\hat{\boldsymbol{\theta}}(\varepsilon))\quad\forall{t}\in\mathbb{R}\quad\implies\quad\theta_{\ell}=\hat{\theta}_{\ell}(\varepsilon)\quad\forall\varepsilon\in\mathbb{R},$ (18)
holds for all parameter symmetries whenever $\theta_{\ell}$ is a universal parameter invariant. Take a radius $\rho>0$ and form the ball $\mathcal{B}_{\rho}\left(\boldsymbol{\theta}\right)$. We need to prove that there exists a radius $\rho$ such that the parameter symmetries for which the implication in Eq. (18) holds also cover the whole ball $\mathcal{B}_{\rho}\left(\boldsymbol{\theta}\right)$. By “covering the whole ball”, we mean that any $\boldsymbol{\theta}^{\star}\in\mathcal{B}_{\rho}(\boldsymbol{\theta})$ yielding identical outputs for all $t\in\mathbb{R}$ must be related to $\boldsymbol{\theta}$ through a parameter symmetry, i.e., $\boldsymbol{\theta}^{\star}=\hat{\boldsymbol{\theta}}(\varepsilon^{\star})$ for some $\varepsilon^{\star}\in\mathbb{R}$. Infinitesimally, the linearised symmetry conditions in Eq. (11) state that the only directions in parameter space that preserve the observed outputs are indeed spanned by the set of parameter infinitesimals $\boldsymbol{\chi}=(\chi_{1},\ldots,\chi_{p})$ defining all parameter symmetries of the output system. Because $\rho$ may be chosen arbitrarily small, there exists a neighbourhood $\mathcal{B}_{\rho}(\boldsymbol{\theta})$ that is entirely covered by parameter symmetries, which implies local structural identifiability of $\theta_{\ell}$. ∎
@@
We say that a model is locally structurally identifiable if all parameters are locally structurally identifiable. In light of Theo. 1, a model is locally structurally identifiable if all parameters are universal parameter invariants. Given our previous discussion about the number of parameter invariants of parameter symmetries, an equivalent formulation is that a model is locally structurally identifiable if the only parameter symmetry of its input-output system is the trivial parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}:\boldsymbol{\theta}\mapsto\boldsymbol{\theta}$.
@@
When a model is locally structurally non-identifiable, it is of interest to find the structurally identifiable parameter groupings or parameter quantities. Assume that the parameter symmetries $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ of interest have $\tilde{p}\leq{p}$ universal parameter invariants denoted by $I_{k}$, $k\in\{1,\ldots,\tilde{p}\}$ which are collected in a vector $\vec{I}_{\boldsymbol{\theta}}$. Crucially, we can always re-parametrise outputs $\mathbf{y}(t,\boldsymbol{\theta})$ in terms of these universal parameter invariants giving us $\mathbf{y}(t,\vec{I}_{\boldsymbol{\theta}})$, and then apply Theo. 1 on the re-parametrised outputs to give:
@@
Cor. 1 (Local structural identifiability of parameter combinations in terms of universal parameter invariants).
@@
The locally structurally identifiable parameter quantities of the output ODE system in Eq. (4) are given by its universal parameter invariants.
The result in Cor. 1 highlights an important difference between the standard differential algebra approach for determining structural identifiability and the approach based on parameter symmetries. The former approach finds the globally identifiable parameter quantities whereas the latter finds locally structurally identifiable parameter quantities. Moreover, the globally structurally identifiable parameter quantities can always be written as a function of the locally structurally identifiable parameter quantities and, based on Cor. 1, this means that any globally structurally identifiable parameter quantity can indeed be written as a function of the universal parameter invariants. Subsequently, we show that these conclusions are also true in the general case. We begin by providing an explanation for why the standard differential algebra approach always finds universal parameter invariants.
We demonstrate that the the standard differential algebra approach for analysing structural identifiability will always find universal parameter invariants (Theo. 1 and Cor. 1). To this end, consider a system of output ODEs in Eq. (4) where the function $\Delta$ is a vector-valued multivariate polynomial for which the monomials are composed of the outputs $\mathbf{y}$ and derivatives of the outputs. In this case, the standard differential algebra approach is conducted in two steps where we first collect all coefficients of the monomials, and then simplify these coefficients algebraically. This procedure yields all identifiable parameter quantities, and here we show that these identifiable parameter quantities are always given by universal parameter invariants according to our definition of local structural identifiability (Cor. 1). This fact is a consequence of two fundamental properties of differential invariants of Lie symmetries.
[…]
Proof.
Thus, we have
[…]
$=\sum_{\ell=1}^{p}\chi_{\ell}\sum_{j=1}^{\tilde{p}}\frac{\partial I_{j}}{\partial\theta_{\ell}}\frac{\partial F}{\partial I_{j}}$
$=\sum_{\ell=1}^{p}\sum_{j=1}^{\tilde{p}}\chi_{\ell}\frac{\partial I_{j}}{\partial\theta_{\ell}}\frac{\partial F}{\partial I_{j}}$
$=\sum_{j=1}^{\tilde{p}}\frac{\partial F}{\partial I_{j}}\underset{=X_{\boldsymbol{\theta}}(I_{j})}{\underbrace{\left(\sum_{\ell=1}^{p}\chi_{\ell}\frac{\partial I_{j}}{\partial\theta_{\ell}}\right)}}$
$=\sum_{j=1}^{\tilde{p}}\frac{\partial F}{\partial I_{j}}\underset{=0}{\underbrace{X_{\boldsymbol{\theta}}(I_{j})}}=0,$ (20)
which is the desired result. ∎
As a short detour, we clarify the difference between parameter invariants (Defn. 13) and universal parameter invariants (Defn. 7) using this fundamental property of differential invariants. Previously, we saw that the trivial parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}:\boldsymbol{\theta}\mapsto\boldsymbol{\theta}$ is always a parameter symmetry of all possible models, and, more importantly, all parameters $\boldsymbol{\theta}$ are parameter invariants of $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}$. Now, take another parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ that has a parameter invariant $I=F(\boldsymbol{\theta})$ defined by some non-constant, non-linear, multivariate and arbitrary function $G$. Moreover, let us assume that $I=G(\boldsymbol{\theta})$ is a universal parameter invariant. Then, the parameter invariant $I=G(\boldsymbol{\theta})$ of $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ is also a parameter invariant of the trivial parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}$ since $I$ is a function of the parameter invariants $\boldsymbol{\theta}$ of $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}$ (Prop. 1). For the same reason, all parameter invariants of $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ are necessarily parameter invariants of $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}$. Nevertheless, the converse statement is not true as many of the parameter invariants of the trivial symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}$ are not shared with $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$.
Using the fact that any function of invariants is itself an invariant, we draw important conclusions about the structure of the system of output ODEs by analysing the linearised symmetry conditions defining parameter symmetries. By definition, the function $\Delta$ in Eq. (4) defining the output-system of ODEs solves the linearised symmetry conditions in Eq. (11) defining all parameter symmetries. This implies that the function $\Delta$ is, in fact, a universal invariant. As a consequence of Prop. 1, the system of output ODEs can therefore always be written as a function of universal invariants.
Prop. 2 (ODEs as functions of universal invariants).
Consider the system of output ODEs defined by a function $\Delta$ as in Eq. (4). Assume that this system of output ODEs has $\tilde{p}\in\{1,\ldots,p\}$ functionally independent universal parameter invariants, and denote these by $I_{k},k\in\{1,\ldots,\tilde{p}\}$ which are collected in a vector $\vec{I}_{\boldsymbol{\theta}}$. Then, $\Delta$ is a function $\Phi$ of the universal invariants according to
Armed with Prop. 21, we understand why the standard differential algebra approach for conducting a global structural identifiability analysis finds identifiable parameter quantities from a symmetry perspective. Again, consider the case discussed previously when the system of output ODEs in Eq. (4) is defined by a function $\Delta$ that is a vector-valued multivariate polynomial. The standard differential algebra approach for elucidating structural identifiability extracts coefficients of the monomials, and by virtue of Eq. (21) these coefficients must either be a constant or a universal parameter invariant. Better still, our notion of structural identifiability in terms of universal parameter invariants (Theo. 1 and Cor. 1) does not only yield locally identifiable parameter quantities, but it also allows us to characterise the parameter transformations preserving the observed outputs in the form of parameter symmetries.
Thus far, we have introduced a theoretical framework for analysing local structural identifiability through parameter symmetries. We now turn to concrete examples that demonstrate how this parameter-symmetry-based approach can be implemented in practice to assess local structural identifiability.
4 Analysing local structural identifiability using parameter symmetries
Subsequently, we use the toy example presented in Section 1, see Eqs. (1) and (2), to illustrate that calculating universal parameter invariants provides information that is identical with the findings of the standard differential algebra approach when the globally and locally structurally identifiable parameter combinations are the same. Thereafter, we demonstrate through the use of a second example that the conclusions of the approach based on calculating universal parameter invariants and the standard differential algebra approach are different when the globally and locally structurally identifiable parameter combinations are different. We proceed by generalising the approach for analysing local structural identifiability using parameter symmetries implemented in these initial two examples, by presenting a three-step procedure called the CaLinInv recipe. Using this three-step procedure, we analyse the local structural identifiability of two larger models.
4.1 Analysing the structural identifiability of a toy model using parameter symmetries
Consider the toy example in Section 1 that results in the output system in Eq. (2) given by $\mathrm{d}y/\mathrm{d}t=(\kappa_{1}+\kappa_{2})-\lambda{y}$. The standard differential algebra approach concluded that the parameter $\lambda$ is globally structurally identifiable and that the parameters $\kappa_{1}$ and $\kappa_{2}$ are individually non-identifiable while their sum $\kappa_{1}+\kappa_{2}$ is globally structurally identifiable.
@@ 4.1 Analysing the structural identifiability of a toy model using parameter symmetries
As a comparison, we next consider the symmetry-based approach for elucidating structural identifiability which entails finding universal parameter invariants of the output ODE in Eq. (2). To this end, we look for parameter symmetries $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ of the output ODE $\mathrm{d}y/\mathrm{d}t=(\kappa_{1}+\kappa_{2})-\lambda{y}$ with the following structure
$\Gamma_{\varepsilon}^{\boldsymbol{\theta}}:(\kappa_{1},\kappa_{2},\lambda)\mapsto(\hat{\kappa}_{1}(\kappa_{1},\kappa_{2},\lambda,\varepsilon),\hat{\kappa}_{2}(\kappa_{1},\kappa_{2},\lambda,\varepsilon),\hat{\lambda}(\kappa_{1},\kappa_{2},\lambda,\varepsilon)).$ (22)
@@
We denote the generating vector field of the parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ in Eq. (22) by
$X_{\boldsymbol{\theta}}=\chi_{\kappa_{1}}(\kappa_{1},\kappa_{2},\lambda)\partial_{\kappa_{1}}+\chi_{\kappa_{2}}(\kappa_{1},\kappa_{2},\lambda)\partial_{\kappa_{2}}+\chi_{\lambda}(\kappa_{1},\kappa_{2},\lambda)\partial_{\lambda}.$ (23)
@@
Given this vector field, we consider the following linearised symmetry condition:
@@
$X_{\boldsymbol{\theta}}\left[\dfrac{\mathrm{d}y}{\mathrm{d}t}-((\kappa_{1}+\kappa_{2})-\lambda{y})\right]=0\quad\textrm{whenever}\quad\dfrac{\mathrm{d}y}{\mathrm{d}t}-((\kappa_{1}+\kappa_{2})-\lambda{y})=0,$ (24)
@@
which states that the solution manifold is invariant under transformations by the parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ in Eq. (22). Carrying out the differentiation on the left-hand side yields the following equivalent equation:
@@
$X_{\boldsymbol{\theta}}=\chi_{\kappa_{1}}(\kappa_{1},\kappa_{2},\lambda)(\partial_{\kappa_{1}}-\partial_{\kappa_{2}}),$ (27)
@@
$X_{\boldsymbol{\theta}}(I(\kappa_{1},\kappa_{2},\lambda))=\chi_{\kappa_{1}}(\kappa_{1},\kappa_{2},\lambda)\left(\frac{\partial I}{\partial\kappa_{1}}-\frac{\partial I}{\partial\kappa_{2}}\right)=0.$ (28)
@@
where $I_{2}$ is an arbitrary integration constant. Since $I_{2}$ is a first integral of Eq. (34), this implies that it is also another differential invariant. In total, this implies that the two universal parameter invariants of the parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ in Eq. (22) are given by
@@
$\Gamma_{\varepsilon}^{\boldsymbol{\theta}}:(\kappa_{1},\kappa_{2},\lambda)\mapsto(\kappa_{1}+\varepsilon,\kappa_{2}-\varepsilon,\lambda),$ (37)